OpenStudy (anonymous):

Solve using the Quadratic Formula 5x^2+7x-2=4x-1

OpenStudy (mertsj):

Get it equal to 0 and then plug into the formula.

OpenStudy (anonymous):

5x^2+3x-1=0

OpenStudy (mertsj):

Now a = 5, b = 3, c = -1 Plug into the formula.

OpenStudy (anonymous):

ahh i seee

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

@drotzmann The denominator should be 2a

OpenStudy (anonymous):

yeah... i meant to write that i didnt realize i for got that sorry haha

OpenStudy (anonymous):

First : We have : \[5x^2+7x-2=4x-1\iff 5x^2+7x-2-4x+1=0\\\iff 5x^2+3x-1=0\\\] So : \[a=5~~ b=3~~c=-1\] Then \[\Delta =b^2-4ac=9+4\times5\times 1=54>0\] There exists 2 solutions : \[x_1=\frac{-b+\sqrt\Delta}{2a}=\frac{-3+\sqrt{54}}{10}\\x_2=\frac{-b+\sqrt\Delta}{2a}=\frac{-3-\sqrt{54}}{10}\]

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\] \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] there i fixed it

OpenStudy (anonymous):

\[\frac{ -3\pm \sqrt{29} }{ 6 }\] the answer or you can simplify it even more

OpenStudy (anonymous):

(-2.102, -3.897) these should be the two answers